OpenStudy (anonymous):

The November 16, 2009 issue of the New York Times reported that there were 671,616 foreign students enrolled at US colleges and universities in 2008-09 — representing an 8% increase from the previous year. It was the largest percentage increase in more than 25 years.

OpenStudy (anonymous):

How many foreign students were studying in US colleges and universities during 2007-2008?

OpenStudy (anonymous):

x+ 8% of x = 671,616

OpenStudy (anonymous):

Solve for x ... CAN U?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

http://openstudy.com/code-of-conduct for more information. Thanks.

OpenStudy (anonymous):

... uh ok?

OpenStudy (nubeer):

suppose there was x students in 2007-2008 so 8% increase , 8/100 (x) 0.08x so x+0.08x = 671,616 can you find x?

OpenStudy (anonymous):

i did but the answer came out totally different

OpenStudy (nubeer):

what is the correct answer?

OpenStudy (anonymous):

621,867 foreign students

OpenStudy (anonymous):

but i got 617886.72

OpenStudy (nubeer):

hmm maybe solve again.. i am getting the right answer with the equation i wrote.

OpenStudy (anonymous):

for which number

OpenStudy (anonymous):

u got 621867?

OpenStudy (nubeer):

x+0.08x =671,616 1.08x=671,616 divide both sides by 1.08

OpenStudy (anonymous):

wth..... i just got the answer but yyyy would u divide i dont get it

OpenStudy (nubeer):

bcos u have to find the x.. x is the number of students in 2007

OpenStudy (nubeer):

its 1.08x right now.. so to make x alone or isolate divide both sides by 1.08

OpenStudy (anonymous):

ohhhh ok so its y=p(1+r)

OpenStudy (nubeer):

oh mm what is this lol?

OpenStudy (anonymous):

thats the formula for the problem i used 671616=p(1+.08)

OpenStudy (nubeer):

ohh ok .. well i am not familiar with that formula.. but ya ok if that formula works fine then its ok

OpenStudy (anonymous):

thank u

OpenStudy (nubeer):

you are welcome.